wiki:UC20190903

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 9/03/19

Attending: Mark Klaene (APO), Kevin Schlaufman (JHU), Ben Williams (UW), Mukremin Kilic (OU), Anne Verbiscer (UVa), Joanne Hughes Clark (Seattle U), John Bally (CU), Sarah Tuttle (UW), Nancy Chanover (NMSU), Leslie Hebb (NAPG), Derek Buzasi (FGCU), Kal Kadlec (UW)


User feedback and comments from institutional representatives:

  • JHU - nothing to report
  • Wyoming - no report
  • Colorado - nothing to report
  • Georgia State - no report
  • Washington - nothing to report
  • UVa - nothing to report
  • FGCU - happy to have finally gotten some data!
  • Seattle - nothing to report
  • NAPG -nothing to report
  • Oklahoma -nothing to report
  • NMSU - question from a user about if/when ARCTIC CCDs will be replaced given the substantial dark current issues. John Bally reported that he takes long exposures in narrow-band filters, for which he needs good dark frames to remove the image structure since the noise is pretty substantial. He has experience in doing this and would be happy to talk with the NMSU user about his approach for mitigating the problem.

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 8/01/2019 - 8/29/2019

0) Overview

Post shutdown observing continued and despite some expected loss time due to weather we continued to be able to observe much of the time. We had minimal problems this period. We adjusted altitude preload as we found a drift after the post shutdown adjustment. DIS was having some readout issue which we believe to be mostly due to high daytime ambient temperatures. Most of the engineering staff time was devoted to KOSMOS II.

1) Telescope

Nominal operations with minimal problems.

2) Instruments

DIS: The DIS blue camera continues to see contamination issues while red camera has stabilized and may start seeing improvement with cooling temperatures. There are no plans to service DIS Blue at this time.

TripleSpec is operational.

Agile is operational.

Echelle is operational. The IOL is steadily getting worse , we are monitoring it closely to determine when and IF we will intervene. The blue value has gotten worse than our nominal limit and red is stable. We expect blue will recover some with cooler temperatures so we are watching it regularly. Let us know if it will effect your science.

NIC-FPS is operational.

ARCTIC is operational; we are working on some minor filter issues.

3) ARCSAT

Operating nominally. The new guide scope mount is in fabrication. The visible sky camera is undergoing a makeover.


Additional telescope and instrument discussion:

Operations have been good other than a few minor tweaks. The scattered light on DIS red has been holding steady for some time, suggesting that we may have gotten through the summer without it getting worse. However, DIS blue is getting worse. Every time we have serviced it during the past 1.5 yrs or so it only lasts for a short period of time. The root cause of the problem was identified as issues with the dewar design, but this would require substantial effort and financial resources to invoke a long term fix. We are no longer routinely servicing DIS blue because it only lasts about a week. That said, if we have users who need DIS blue for their science and we can service it around the time of their scheduled observations we are not adverse to doing so, although they would need to scheduled to observe with it almost immediately after the servicing. The inter-order light on the echelle has gotten a little worse (seen more on the blue side); we are monitoring it closely and we think/hope it may get better as the temperatures start to cool down.


Q3 3.5m and ARCSAT scheduling:

There are currently 3 open halves left in Q3; please submit a request if needed. We will possibly be moving some requested Q4 time into Q3 because October A halves are fairly oversubscribed and we now have some room in Q3 since KOSMOS commissioning has been postponed (see below). We still have some open time on ARCSAT in September. Please submit requests for any open Q3 time to Nancy, Ben, Russet, and cc your institutional scheduler.


Q4 3.5m and ARCSAT scheduling:

We now have all requests in and Russet is working on the schedule. There are lots of class trips coming to the site in Q4. As a reminder, it helps to know FAR AHEAD OF TIME if anyone plans on bringing a class trip. The schedule should be ready some time next week. The call for Q4 time on ARCSAT will come out after 3.5m Q4 schedule is released. As a reminder to UC reps (please communicate this to the institutional schedulers as well): tell users to request special protection from preemption if they need it!


NA2 Baffle Project:

Work is proceeding on schedule. John Wilson has been in regular contact with Bill and the folks at Photon Engineering, who determined that the stray light in the system is coming from scattering from the optics themselves. This is a good situation to be in because it means that no other aspects the system are contributing to the stray light. Once the design of the new baffle is narrowed down fabrication will begin. We expect it to be installed in early Q1 2020.


KOSMOS Update:

Due to delays with getting several parts finished the KOSMOS commissioning is being pushed to the week of December 8. There is now a clear schedule for the various steps that need to happen leading up to commissioning. The APO engineering staff is rewiring the CCD. The UW TEG will have all parts in house soon and then will begin testing slit viewer and installing the new nanofab slits. A new mechanical engineer, Rishi, was recently hired into the TEG; he is already helping to bring KOSMOS online. They can install any new grisms if they are procured in time, but the nominal plan is to deploy the instrument with what we received from KPNO. The option still exists to include different grisms and slits (we will have a lot of slits to choose from). There is a slit wheel as well as a grism wheel, so these will both be easy to change over the course of a night. With a higher resolution grism the tradeoff would be lower throughput and a narrower bandpass overall. We do expect to be able to make some direct comparisons between KOSMOS and DIS; we expect considerable improvement in throughput over DIS. The cost of new grisms is not certain (ballpark ~ $10k); there has been a shift in the VPH market such that there are probably 2 viable vendors these days, only one of which is responsive to astronomy applications.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • Getting DAWG underway: the group had their first telecon and is gearing up to survey users concerning available software for data analysis. Next telecon will be at the end of September.
  • Getting narrow-band filters for ARCTIC: Nancy sent out a survey to 35m-general soliciting input concerning desired narrow-band filters that can take advantage of the full (soon-to-be-unvignetted) FOV of ARCTIC, as well as additional desired gratings for KOSMOS. UC reps were asked to remind their users to complete the survey by the deadline of September 20.
  • Science communications intern: Nancy is preparing to hire an undergraduate from NMSU's Journalism and Mass Communications department as a science communications intern. The position will be posted in September.
  • AAS presence: Nancy will reserve an exhibitor booth, which will be used for small-group meetings of APO-affiliated meeting attendees and for promotional purposes (new member recruitment, collaboration building, etc.). UC reps were asked to poll their users to see who is going to the AAS meeting and would be willing to help out with booth presence.
  • Student training videos: There are several class trips scheduled for this fall (more than 4 in Q4!), so Nancy would like to revisit the idea of having the students do a video hackathon (students filming themselves) to create some training videos. After some brainstorming on the telecon we came up with some initial topics to try out:
    • enter coordinates and slew to a target, find a guide star, and set up the guiding
    • examining different operational scenarios for guiding (e.g. what if guide star is bright, what if it's too faint, what if there are none? what if it's cloudy so the guide star fades in and out?)
    • how to collect calibration data for each instrument
    • loading target lists into TUI (format of file, how to load it)
    • things that make APO unique from other observatories (instrument change procedure, i.e. which ones don’t need it, NA2 port vs others, properties of the NA2 guider, the echelle front end and how to use calibration and flat field lamps for the echelle)

Nancy will contact trip leaders to see what might be possible.

  • Bill reached out to Steward Observatory to inquire about a) their portable mirror coating capability and b) the turn-key 6.5m telescope concept that they are developing. Three individuals from Steward, M3 Engineering and CAID Industries came to APO on August 22 to tour the site to examine the available infrastructure and give their sales pitch about the telescope. The cost is under $70M. If we could find a daytime use partner (e.g. military) then that could be reduced substantially. We would need to evaluate whether or not ARC partners (current and/or future) would be interested in pursuing this further; this is clearly a topic that will require more discussion at future strategic planning meetings.

Request for Science Highlights for BoG report

Nancy is requesting science highlights from each ARC member institution and leasing partner to present at the ARC Board of Governors meeting in early November. The deadline for submitting them is October 15. UC reps are asked to collect these from their users and send them to Nancy or have them sent directly. There is no template to follow but the slides should include project title, PI and team member names, institution, a graphic, some bullet points summarizing the take-away result, and journal references (if available).


ACTION ITEMS:

Open action items from previous meetings:

  • Ben will send out a link to KOSMOS info for interested prospective users. STATUS: OPEN. This is intended to be a page where prospective users could learn about the general capabilities of KOSMOS at a glance. Sarah and Kal have been worrying more about getting KOSMOS ready for commissioning, so this remains open. Sarah envisions this to be an entry point with capabilities that will be built up to a full manual that they can link to the APO web site. Ben will work with Kal to give him examples of what we're looking for. STATUS: OPEN. The document is pretty much ready to go; Ben will contact Nancy and Mark offline to determine where/how to post it.
  • UC reps should notify users of any Q3 OPEN/DD time remaining. STATUS: OPEN.
  • Nancy will coordinate with Mike Crenshaw regarding distribution of the Astro2020 APC White Paper. STATUS: OPEN
  • Nancy and Misty will survey the 3.5m user community regarding any potential additional grisms that would be desired for KOSMOS. STATUS: OPEN
  • Nancy will identify some possible topics for training video hacks to be done by visiting classes of students. STATUS: OPEN

New action items from this meeting:

  • UC reps should remind their users to complete the survey about desired narrow-band ARCTIC filters and KOSMOS gratings by the deadline of September 20.
  • UC reps should poll their users to see who is going to the AAS meeting and would be willing to help out with booth presence.
  • Nancy will contact trip leaders bringing student groups to APO in Q4 to discuss training video ideas.
  • UC reps should request science highlights from their users and have them sent to Nancy no later than October 15.

All Other Business

None.


Next meeting:

The next Users Committee telecon will be on Monday October 7 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.